OpenStudy (anonymous):

graph the equation-3x + 5y=10

OpenStudy (m):

easiest way is to plot points. solve for x and solve for y. plug in numbers such as 1, 2 ,3, 4 etc. and plot the results in the xy graph.

OpenStudy (anonymous):

5y = 3x +10 y = (3x + 10)/5 now you can choose several numbers to be your x, plug them is the equation to find y, mark the points on your graph and draw the function PS: your x can be any number for example 1, 2, -2, 3.5 and so on.

OpenStudy (anonymous):

(0,2)(3,-1) CORRECT ME IF WRONG

OpenStudy (anonymous):

M=3/5 B=2